### v2

# through $\mathcal{R}$eal-$\mathcal{T}$ime $\mathcal{V}$ideo

Abstract
Multimodal Large Language Models (MLLMs) increasingly excel at perception, understanding, and reasoning. However, current benchmarks inadequately evaluate their ability to perform these tasks continuously in dynamic, real-world environments. To bridge this gap, we introduce $\mathcal{RTV}\text{-}Bench$, a fine-grained benchmark for MLLM real-time video analysis. $\mathcal{RTV}\text{-}Bench$ includes three key principles: (1) Multi-Timestamp Question Answering (MTQA), where answers evolve with scene changes; (2) Hierarchical Question Structure, combining basic and advanced queries; and (3) Multi-dimensional Evaluation, assessing the ability of continuous perception, understanding, and reasoning. $\mathcal{RTV}\text{-}Bench$ contains 552 diverse videos (167.2 hours) and 4,631 high-quality QA pairs. We evaluated leading MLLMs, including proprietary (GPT-4o, Gemini 2.0), open-source offline (Qwen2.5-VL, VideoLLaMA3), and open-source real-time (VITA-1.5, InternLM-XComposer2.5-OmniLive) models. Experiment results show open-source real-time models largely outperform offline ones but still trail top proprietary models. Our analysis also reveals that larger model size or higher frame sampling rates do not significantly boost $\mathcal{RTV}\text{-}Bench$ performance, sometimes causing slight decreases. This underscores the need for better model architectures optimized for video stream processing and long sequences to advance real-time video analysis with MLLMs. Our benchmark toolkit is available at Project Page.
**footnotetext: Both authors made equal core contributions.$\dagger$\$$\dagger$\$footnotetext: Corresponding author.
Figure 1: Specific examples of different tasks in RTV-Bench.
and
represent the fundamental questions in the question group, and the answers are underlined.
is a dynamically answered question, and the correct answer is determined by the time of the query. The correct answer to the query will change at different times, and we have marked the correct answers that should be responded to at different query times.
1 Introduction
The ability to comprehend and respond to complex real-world scenarios in real-time represents a crucial challenge in the pursuit of general artificial intelligence Dang et al. (2024); Zhang et al. (2025b); Hu et al. (2025). Drawing inspiration from the remarkable achievements of large language models (LLMs) in a wide range of tasks Liu et al. (2024); Huang et al. (2025), the potential of multimodal large language models (MLLMs) is increasingly being realized in the domain of visual scene analysis. The research trajectory of Video-LLMs Lin et al. (2023); Maaz et al. (2024); Cheng et al. (2024); Wang et al. (2024a) is progressively advancing from the understanding of short, vision-centric videos Li et al. (2023a); Jin et al. (2024); Lin et al. (2024a); Li et al. (2024d) to the integration of omni-modal information—including video, audio, and subtitles Chen et al. (2023); Duan et al. (2023); Li et al. (2023b); Tan et al. (2024); Li et al. (2024c)—for comprehensive long video analysis.
Recently, VStream Zhang et al. (2024a) was the first to attempt to test this capability, with a focus primarily on extending the duration of videos. In addition, both StreamingBench Xiong et al. (2025) and OVOBench Li et al. (2025) have made varying degrees of improvements in the types and standards of assessment. However, their evaluation of real-time responsiveness is often inadequate, overlooking the capacity to capture transitions and fleeting details from visual input that arrive sequentially, not instantaneously. This limitation highlights the need for a more focused assessment of continuous analysis abilities.
Based on the above considerations, we introduce $\mathcal{RTV}\text{-}Bench$, featuring three core innovations designed to benchmark the continuous analysis capabilities—specifically perception, understanding, and reasoning—of MLLMs within real-time video contexts. First, the Multi-Timestamp Q&A mechanism challenges real-time tracking and state update by posing queries whose answers evolve within a video. Crucially, unlike benchmarks like OVO-Bench Li et al. (2025) that typically introduce different questions at different timestamps, RTV-Bench revisits the same conceptual query, where only the correct answer shifts as the scene unfolds. This approach more rigorously tests the model capacities for continuous analysis in real-time scenarios, surpassing single-query-single-answer and static evaluations. Second, the Hierarchical Question Structure enforces reliable, sequential reasoning by employing a basic-to-advanced design where higher-order questions logically depend on grasping foundational perceptions and understanding, thus mitigating cognitive shortcuts. Finally, Multidimensional Evaluation moves beyond single scores to provide granular diagnostics, assessing performance across eight fine-grained dimensions critical to continuous analysis in dynamic scenarios, offering a more insightful perspective on model capabilities and bottlenecks in handling real-time video information, detailed cases are shown in Figure 1.
Through RTV-Bench’s novel design and the detailed evaluations presented herein, we provide comprehensive insights into the current state of MLLMs for continuous video analysis: Results for all tested models on RTV-Bench reveal significant bottlenecks for existing MLLMs in real-time video analysis, as (1) most models achieve accuracy below 50%. Furthermore, (2) model performance shows minimal correlation with the number of input frames or model size. Notably, (3) models specifically designed for real-time video significantly outperform traditional offline video models; even the lowest-performing real-time model tested (VITA-1.5 Fu et al. (2025)) surpasses a traditional offline counterpart (VideoLLaMA2 Cheng et al. (2024)). Building upon these findings, we further discuss promising research trajectories to enhance MLLM online video processing. We expect that our benchmark, analysis, and outlined future directions will contribute significantly to the development of more robust and capable online video MLLMs.
2 RTV-Bench
Figure 2: Video Categories and Distribution of Question Difficulty and Query Characteristics. (Left) RTV-Bench overs 3 key domains and 16 sub-class video types. (Center) Distribution of question difficulty levels across eight representative task types, measured by percentage-based performance ranges. (Right) Distribution of question queries by video length, categorized into Shallow, Moderate, and Deep levels. The bar heights indicate counts, while the line chart overlays query proportions for each duration bucket.
In this section, we discuss a common challenge in both long video comprehension and real-time video analysis—the continuous analysis capability of MLLMs. To tackle this, we have meticulously curated a benchmark to assess model performance in real-time and continuous scenarios, and developed new metrics to characterize the performance of multimodal large models in this bench.
2.1 Challenge for Video Analysis
Existing benchmarks for long video analysis Fu et al. (2024); Wu et al. (2025) attempt to gauge the true capacity of MLLMs by extending video duration, enhancing the difficulty of QA tasks, and introducing new sub-tasks. Other works Lin et al. (2024b) focus on real-time video scenarios, evaluating model responsiveness to queries in real-time contexts. However, current models face the risk of memory loss and attention shift, both in long and real-time video contexts. We categorize these challenges as failures in the continuous analysis capability of MLLMs. In video scenarios, MLLMs are considered to possess strong continuous analysis ability only if they can effectively recall prior visual information to accurately respond to user queries (similar to needle-in-a-haystack tasks in natural language settings) and maintain robust perception of incoming streams of new visual data and queries (akin to multimodal multi-turn dialogues).
2.2 Benchmark Overview
The RTV-Bench is designed to assess a model’s ability to perceive, understand, and reason in long video and real-time streaming contexts. Models should be adept at recognizing correct temporal query patterns. For example, in an ongoing sports match, when a user inquires about the goalkeeper’s actions, the model should respond with the situation at the query time, rather than historical moments. However, it should also retrieve relevant information from memory, such as identifying the goalkeeper from earlier footage. This necessitates that models not only possess exceptionally high fundamental scene comprehension capabilities, but also demonstrate effective understanding of all stages within the temporal flow—namely, comprehensive spatiotemporal comprehension encompassing the past, present, and future. To this end, we propose a series of sub-tasks for video continuous analysis, addressing both spatiotemporal elements within the video and the intrinsic capabilities of the model. The intrinsic model capabilities are divided into eight categories: Temporal Perception (TP), Scene Perception (SP), Visual Perception (VP), Future Prediction (FP), Phenomenological Understanding (PU), Intent Analysis (IA), Global Understanding (GU), and Spatiotemporal Reasoning (SR), see Figure 1 for details.
2.3 Benchmark Construction
Terminology
The RTV-Bench dataset comprises $N=552$ videos sourced from the internet. A key characteristic is its question structure designed to directly probe temporal dynamics. For each video, questions are organized into sets. Each question set contains a minimum of three multiple-choice questions.
Crucially, the core design principle revolves around time-varying correct answers (Multi-Timestamp QA): within a set of questions, the same underlying conceptual query (for example, ’What is person A holding?’, ’Where is the car heading?’) is effectively evaluated at different points in time within the video. This means that the correct answer option changes depending on the specific timestamp or time interval referenced or implied by the question context or its associated earliest inferable timestamp (Figure 1). This challenges models not just to locate information, but to actively track changes and update their understanding of a situation as it evolves. This design directly targets the assessment of continuous understanding and the model’s sensitivity to temporal changes, see Figure 4 for cases.
The question set features three questions of escalating difficulty. The first two are simpler, while the final question is significantly more complex, demanding integration of broader context or clues. The information and reasoning needed for the final question generally encompass those required for the first two. Successfully answering the complex third question strongly suggests the capability to also answer the simpler ones. This structure evaluates the model’s ability to handle increasing complexity, synthesize comprehensive context, and perform robust analysis.
In addition, we emphasize diversity through the richness of specific sub-scenes and the distribution of video lengths. The RTV-Bench primarily encompasses intelligent driving, sports events, and egocentric videos—categories rich in dynamic information and real-time contextual relevance. Each category includes a variety of sub-categories, as illustrated in Figure 2.
Benchmark MT Num. of QA Avg. Dur. (min) Total Dur. (h) Labels
Video-MME Fu et al. (2024) 2.7k 17 254.0 M
MSRVTT Xu et al. (2016) 73k 0.25 12.5 A
MSVD-QA Chen and Dolan (2011) 13k 0.17 1.4 A
MovieChat-1k Song et al. (2024) 13k 9.4 156.0 M
MVBench Li et al. (2024b) 4k 0.25 34.5 A+M
ActivityNet-QA Yu et al. (2019) 58k 1.87 25.0 M
Vstream-Q Zhang et al. (2024a) 3.5k 40 21.0
StreamBench Wu et al. (2024) 1.8k 4.5 25.0 M
OVO-Bench Li et al. (2025) 2.8k 6 66.0 A+M
RTV-Bench (Ours) 4.6k 18 167.2 M
Table 1: Comparison of video QA benchmarks. MT: multi-timestamp questions. Labels: A (automatic), M (manual), A+M (both).
Data Statistic
RTV-Bench comprises 552 videos with a total duration of 167.2 hours (average 18.2 minutes per video) and contains 4,631 QA pairs, for detailed information, refer to Figure  2 and Table  1. The benchmark features diverse question scenarios and evenly distributed video durations, thereby establishing a comprehensive framework for video continuous understanding tasks with rich sub-scenarios and structured problem dimensions.
Video Collection and Filtering
Our video data sources include EgoSchema Mangalam et al. (2023) and publicly available online videos. Unlike most existing benchmarks, we incorporated manual review during the collection phase. Three data collectors sourced videos from various domains and manually excluded highly similar videos, focusing on long videos with high dynamics and real-time needs. These sources ensure targeted and scientific evaluation of video models’ real-time capabilities.
Manual Annotation
Rigorous manual annotation by qualified experts was essential for RTV-Bench’s reliability in assessing continuous video analysis. While an LLM (Deepseek Liu et al. (2024)) provided initial question templates, human annotators significantly refined questions to target dynamic scenes and temporal reasoning. They meticulously crafted questions where answers change over time and identified the earliest valid timestamp for each answer option in MTQA. This human-centric, multi-annotator process ensured robustness and created evaluations specifically focused on temporal dynamics in video.
Quality Control
To ensure the benchmark’s quality, each video and its corresponding Q&A pairs underwent multiple rounds of review. We manually filtered videos based on length distribution and the presence of sub-scenes examining real-time event changes, resulting in high-quality videos focused on real-time analysis tasks. We conducted manual video-question alignment and precise timestamp checks, utilizing GPT-4 and human review to verify annotation format, translation, and sensitive information.
3 Experiments
3.1 Experiment Setup
Our experiments were conducted across two A800 GPUs to comprehensively evaluate the performance of mainstream MLLMs on our benchmark. The selected models include VideoLLaMA2 Cheng et al. (2024), VideoLLaMA3 Zhang et al. (2025a), GPT-4O Hurst et al. (2024), InternLM-XComposer2.5-OmniLive (IXC2.5-OL) Zhang et al. (2024b), VITA1.5 Fu et al. (2025), LLaVA-Video Lin et al. (2023), LLaVA-OneVision Li et al. (2024a), and Qwen2.5-VL Yang et al. (2024). These models were primarily tested with 7 billion parameters (7B) to ensure a fair comparison under similar computational constraints.
Real-Time Video Model vs. Offline Video Model
We compare two model categories: traditional offline models ($M_{offline}$) and novel real-time online models ($M_{online}$). These categories differ significantly in architecture, training, and data requirements. Architecturally, $M_{offline}$ typically employ sequential vision encoder-decoders, often limited by fixed context windows and incurring higher processing latency. In contrast, $M_{online}$ are designed for continuous, low-latency ingestion of the video stream $V$. They prioritize maintaining an internal state $S_{t}$ that summarizes the video information processed up to time $t$ (i.e., $V[0,t]$), enabling real-time responsiveness. Models like IXC2.5-OL Zhang et al. (2024b) implement this using techniques like modular parallelism and dedicated long-term memory. These architectural distinctions lead to different training paradigms and data needs. $M_{offline}$ usually rely on end-to-end fine-tuning using standard annotated video datasets. $M_{online}$, however, often require specialized training strategies (e.g., VITA-1.5’s staged fusion, IXC2.5-OL’s targeted training for memory and interaction) and benefit most from specialized corpora designed for long-duration, interactive streaming scenarios. The ability of $M_{online}$ to maintain and update state $S_{t}$ is particularly relevant for RTV-Bench’s Multi-Timestamp QA (MTQA) challenge, where the correct answer $A^{*}(Q,t_{q})$ to a query $Q$ depends on the specific query time $t_{q}$. To evaluate both model types on this benchmark, we adapt the testing procedure. For $M_{online}$, queries $Q$ are presented at their timestamp $t_{q}$, and the models leverage their continuously updated state $S_{t_{q}}$ to generate the answer $A_{M_{online}}=M_{online}(Q,t_{q}|S_{t_{q}})$. Since $M_{offline}$ lack this inherent streaming capability, we simulate real-time interaction for them: when a query $Q_{i}$ is posed at time $t_{q,i}$, we extract and provide only the relevant video segment $V_{i}$ (corresponding to the query’s context). The offline model’s answer is thus based solely on this isolated segment: $A_{M_{offline},i}=M_{offline}(Q_{i},V_{i})$.
To evaluate the performance of these models, we employed two metrics: Accuracy and Score.
Accuracy.
The accuracy metric measures the proportion of correct answers provided by the model compared to the ground truth.
Score.
The score metric evaluates the model’s ability to correctly answer advanced-level questions (type q2), contingent upon its demonstrated mastery of prerequisite basic questions (types q0 and q1) within the same question group, thus emphasizing reliable advanced reasoning built upon a solid foundation. Calculation involves a prerequisite check for each group $i$: if all basic questions are correct ($B_{i}=1$), the group contributes points equal to the number of correctly answered q2 questions ($N_{q2,i}^{\text{correct}}$); otherwise ($B_{i}=0$), it contributes zero points. The final score is the ratio of total conditionally awarded points to the total number of q2 questions across all $N$ valid groups (those containing q2). Formula:
$\text{Score}=\frac{\sum_{i=1}^{N}B_{i}\cdot N_{q2,i}^{\text{correct}}}{\sum_{i=1}^{N}N_{q2,i}^{\text{total}}}$
where $N$ is the number of valid groups, $B_{i}$ is the prerequisite indicator (1 if basics are correct, 0 otherwise) for group $i$, $N_{q2,i}^{\text{correct}}$ is the count of correct q2 answers in group $i$, and $N_{q2,i}^{\text{total}}$ is the total count of q2 questions in group $i$. Advantages of this metric include: ensuring foundational accuracy by rewarding advanced correctness only when basics are mastered; reflecting model robustness by penalizing superficial success on complex tasks without fundamental understanding; and aligning with hierarchical learning principles where complex skills build upon simpler ones.
Model #Size Perception Understanding Reasoning FQA MTQA Overall
Acc (%)/Score Acc (%)/Score Acc (%)/Score Acc (%) Acc (%) Acc (%)/Score
Open-Source Offline Video Models
Qwen2.5-VL 7B 33.53/3.53 31.39/3.70 33.65/5.20 31.37 34.37 32.75/4.10
VideoLLaMA2 7B 40.62/8.67 39.85/7.77 37.49/6.75 45.77 34.95 39.55/7.90
VideoLLaMA3 7B 37.98/5.83 35.29/5.73 35.78/6.80 38.62 34.91 36.42/6.10
LLaVA-OneVision 7B 35.38/3.97 34.21/4.63 33.57/4.95 35.80 33.58 34.49/4.40
LLaVA-Video 7B 35.83/5.03 33.81/3.77 35.15/5.75 36.28 34.17 34.90/4.80
Open-Source Online Models
VITA-1.5 7B 45.66/12.80 44.12/11.83 43.37/10.15 55.06 36.32 44.51/11.80
IXC2.5-OL 7B 47.21/15.87 48.22/15.23 46.18/14.45 59.05 38.21 47.33/15.40
Closed-Source Business Models
GPT-4o 51.61/21.90 49.31/20.76 48.71/23.95 56.53 44.73 50.02/22.10
Gemini 2.0 Flash 41.67/11.00 42.71/12.73 41.44/12.05 47.49 38.64 42.00/12.00
Table 2: Evaluation results on RTV-Bench. Perception, Understanding and Reasoning are different tasks. FQA notes that Foundational Video Question&Answersing without multi-timestamp answers. MTQA notes that Multi-Timestamp Question&Answersing with multi-timestamp answers.
Method #Size TP VP SP Overall
Acc (%) / Score Acc (%) / Score Acc (%) / Score Acc (%) / Score
Open-Source Offline Video Models
Qwen2.5-VL 7B 32.37 / 2.4 37.48 / 5.8 30.73 / 2.4 33.53 / 3.53
VideoLLaMA2 7B 39.52 / 7.9 42.49 / 9.4 39.85 / 8.7 40.62 / 8.67
VideoLLaMA3 7B 37.82 / 6.1 39.24 / 7.6 36.87 / 3.8 37.98 / 5.83
LLaVA-OneVision 7B 35.09 / 3.9 35.86 / 3.8 35.20 / 4.2 35.38 / 3.97
LLaVA-Video 7B 34.07 / 4.8 38.97 / 5.8 34.45 / 4.5 35.83 / 5.03
Open-Source Online Models
VITA-1.5 46.51 / 12.1 47.09 / 13.2 43.39 / 13.1 45.66 / 12.8
IXC2.5-OL 7B 49.57 / 17.6 49.80 / 16.5 42.27 / 13.5 47.21 / 15.87
Closed-Source Business Models
GPT-4o 48.60 / 18.2 53.59 / 23.4 52.63 / 24.1 51.61 / 21.90
Gemini 2.0 Flash 40.49 / 9.5 45.19 / 16.1 39.34 / 7.4 41.67 / 11.00
Table 3: Detailed evaluation results on the category of Perception. Temporal Perception (TP), Visual Perception (VP) and Scene Perception (SP)
Method #Size GU PU IA Overall
Acc (%) / Score Acc (%) / Score Acc (%) / Score Acc (%) / Score
Open-Source Offline Video Models
Qwen2.5-VL 7B 29.11 / 2.2 35.69 / 6.9 29.36 / 2.0 31.39 / 3.70
VideoLLaMA2 7B 37.34 / 7.6 42.21 / 9.5 40.92 / 6.2 39.85 / 7.77
VideoLLaMA3 7B 33.54 / 5.8 39.13 / 5.9 33.39 / 4.3 35.35 / 5.33
LLaVA-OneVision 7B 32.07 / 4.3 33.51 / 3.0 37.06 / 6.6 34.21 / 4.63
LLaVA-Video 7B 29.42 / 2.5 35.69 / 3.9 36.33 / 4.9 33.81 / 3.77
Open-Source Online Models
VITA-1.5 7B 40.30 / 7.2 46.01 / 15.1 46.06 / 13.2 44.12 / 11.83
IXC2.5-OL 7B 43.88 / 11.9 52.17 / 18.7 48.62 / 15.1 48.22 / 15.23
Closed-Source Business Models
GPT-4o 45.02 / 15.7 54.32 / 25.8 48.58 / 20.8 49.31 / 20.76
Gemini 2.0 Flash 35.70 / 10.6 45.65 / 11.3 46.78 / 16.3 42.71 / 12.73
Table 4: Detailed evaluation results on the category of Understanding. Phenomenological Understanding (PU), Global Understanding (GU) and Intent Analysis (IA)
Method #Size FP SR Overall
Acc (%) / Score Acc (%) / Score Acc (%) / Score
Open-Source Offline Video Models
Qwen2.5-VL 7B 35.33 / 4.4 33.67 / 4.5 34.50 / 4.45
VideoLLaMA2 7B 41.47 / 7.5 33.50 / 6.0 37.49 / 6.75
VideoLLaMA3 7B 38.05 / 6.9 33.84 / 3.9 35.95 / 5.40
LLaVA-OneVision 7B 38.23 / 7.2 28.91 / 2.7 33.57 / 4.95
LLaVA-Video 7B 39.08 / 9.1 31.22 / 2.4 35.15 / 5.75
Open-Source Online Models
VITA-1.5 7B 47.95 / 12.2 38.78 / 8.1 43.37 / 10.15
IXC2.5-OL 7B 51.88 / 18.1 40.48 / 10.8 46.18 / 14.45
Closed-Source Business Models
GPT-4o 54.67 / 27.1 42.75 / 20.8 48.71 / 23.95
Gemini 2.0 Flash 44.42 / 13.6 38.46 / 10.5 41.44 / 12.05
Table 5: Detailed evaluation results on the category of Reasoning. Future Prediction (FP) and Spatiotemporal Reasoning (SR).
3.2 Experiment Results
Online vs. Offline Models.
RTV-Bench results (shown in Table 2) demonstrate that online models designed for real-time processing, particularly IXC2.5-OL, outperform offline models in overall performance (e.g., IXC2.5-OL: 47.33% Acc / 15.40 Score vs. VideoLLaMA2: 39.55% Acc / 7.90 Score). This advantage is notable in tasks requiring temporal analysis, such as Multi-Timestamp Question Answering (MTQA), where online models like IXC2.5-OL (38.21% Acc) surpass their offline counterparts (typically 33-35% Acc). We observe that (Table 4) all models excel at Visual Perception tasks, likely because these tasks are more image-centric compared to Temporal or Scene Perception, and current video models often build upon image architectures. This suggests a need for more specialized video model designs, such as dedicated spatio-temporal architectures or real-time processing capabilities.
Open-Source vs. Close-Source Models.
However, a significant gap remains compared to leading closed-source models like GPT-4o. GPT-4o exhibits superior performance across the board, especially in Overall Score (22.10) and on complex tasks like Understanding and Reasoning (Tables 4 and 5). Evidently, top-tier closed-source models still hold a significant edge, particularly in complex cognitive capabilities.
Accuracy vs. Score.
The Score metric is crucial for revealing reliable, hierarchical reasoning beyond simple Accuracy. While GPT-4o’s Accuracy gain over IXC2.5-OL is moderate (50.02% vs. 47.33%), its substantially higher Score (22.10 vs. 15.40) indicates a more robust reasoning process, reliably building upon foundational understanding to address complex queries. Conversely, lower Scores relative to Accuracy, common among open-source models, suggest instability in multi-step reasoning. Thus, as shown in Figure 3.the Score metric effectively quantifies the deeper, more reliable comprehension capabilities demonstrated by models like GPT-4o on RTV-Bench’s challenging tasks and it can distinguish model performance more clearly.
Impact of Model Scales.
To investigate the effect of model scale, we evaluated Qwen2.5-VL with different parameter sizes. As depicted in Figure 3, the performance across different model sizes exhibits limited variation. Accuracy remains relatively consistent across the tested configurations. For instance, Overall Accuracy hovers narrowly between approximately 31.4% and 32.8%. The limited benefit from scaling suggests the core challenge for real-time performance lies less in model size and more in model designs. Larger models failing to improve significantly implies a potential bottleneck in how effectively their internal representations capture the demands of real-time processing.
Impact of Frame Numbers.
Figure 3 presents another counter-intuitive finding. While more input frames typically yield better performance in offline evaluations, our real-time setting shows no performance improvement with increased frames. In fact, IXC2.5-OL exhibits significant degradation. This suggests potential limitations in the model’s processing capacity or a breakdown mechanism, such as attention dilution. A potential limitation is that the current frame numbers may not adequately capture all critical moments or key information within long videos. Even with increased sampling density, crucial details might still be missed.
Figure 3: Performance visualization and analysis on RTV-Bench: (a) Visualization of overall Accuracy results; (b) Visualization of overall Score results; (c) Performance impact of varying input frame counts; (d) Performance comparison across different Qwen2.5-VL model scales
Figure 4: Comparison of answers from different models for the same question on the same video. Green indicates correct answers or timestamps; red indicates incorrect answers or timestamps. This case demonstrates that even current high-performance models (like Qwen2.5-VL Wang et al. (2024a), Gemini 2.0 Team et al. (2024), and GPT-4o Hurst et al. (2024)) struggle to provide high-quality responses (where both the answer and the corresponding timestamp are accurate).
4 Limitations and Future Work
Our benchmark reveals counter-intuitive findings, such as the limited impact of model scale and input frame count on performance, suggesting that MLLM mechanisms for processing continuous video are poorly understood and effective analysis tools are lacking. Furthermore, the current evaluation is primarily limited to the visual modality. Future work will focus on investigating the underlying causes of these phenomena and developing more adapted analytical methods. Concurrently, a key direction involves incorporating important modalities like audio into RTV-Bench to enable a more comprehensive evaluation of continuous perception, understanding, and reasoning in realistic multimodal scenarios.
5 Related Work
5.1 Advanced MLLMs towards Real-time Video Analysis
Video-oriented large-scale models represent a highly promising domain with significant real-world potential, giving rise to numerous valuable applications Maaz et al. (2024); Fu et al. (2024). Video-LLMs have rapidly advanced, moving from analyzing short videos Li et al. (2023a); Lin et al. (2023) to longer ones Shen et al. (2024); Wang et al. (2024a); Zohar et al. (2024) using various techniques. However, most research targets offline video analysis. Real-world applications necessitate real-time, continuous perception and reasoning on unfolding video streams. While models like InternLM-XComposer2.5-OmniLive Zhang et al. (2024b), VITA-1.5 Fu et al. (2025), and Dispider Qian et al. (2025) are exploring real-time capabilities, rigorously evaluating their ability to continuously track and understand dynamic events remains a critical challenge.
5.2 Existing Video Benchmarks
Existing benchmarks primarily focus on offline evaluation using offline videos Fu et al. (2024); Wang et al. (2024b); Wu et al. (2025). These are less suited for assessing how MLLMs track dynamically changing states, as they often use static question-answer pairs. Newer benchmarks tackle real-time and streaming aspects Lin et al. (2024b); Li et al. (2025), evaluating responsiveness and contextual understanding in online settings. However, RTV-Bench specifically addresses the gap in evaluating continuous perception, understanding, and reasoning. Its key distinction is the use of dynamic question answering, where the correct answer evolves with the video stream, directly probing the MLLM’s ability to maintain and update its understanding of complex, unfolding events over time, complemented by a multi-dimensional evaluation structure.
6 Conclusion
In this work, we introduce RTV-Bench, a Benchmark focuses on evaluating the video continuous analysis capabilities of MLLMs. We proposed relevant evaluation tasks and constructed RTV-Bench, which comprises 552 videos and 4,631 QA pairs. Extensive experiments conducted on diverse models reveal that: Models specifically designed for online video processing exhibit superior video continuous understanding capabilities, whereas simply increasing the number of uniformly sampled frames during training or inference stages currently shows limited effectiveness in enhancing performance on video continuous tasks.