### v3

# through $\mathcal{R}$eal-$\mathcal{T}$ime $\mathcal{V}$ideo

Abstract
Multimodal Large Language Models (MLLMs) increasingly excel at perception, understanding, and reasoning. However, current benchmarks inadequately evaluate their ability to perform these tasks continuously in dynamic, real-world environments. To bridge this gap, we introduce $\mathcal{RTV}\text{-}Bench$, a fine-grained benchmark for MLLM real-time video analysis. $\mathcal{RTV}\text{-}Bench$ includes three key principles: (1) Multi-Timestamp Question Answering (MTQA), where answers evolve with scene changes; (2) Hierarchical Question Structure, combining basic and advanced queries; and (3) Multi-dimensional Evaluation, assessing the ability of continuous perception, understanding, and reasoning. $\mathcal{RTV}\text{-}Bench$ contains 552 diverse videos (167.2 hours) and 4,631 high-quality QA pairs. We evaluated leading MLLMs, including proprietary (GPT-4o, Gemini 2.0), open-source offline (Qwen2.5-VL, VideoLLaMA3), and open-source real-time (VITA-1.5, InternLM-XComposer2.5-OmniLive) models. Experiment results show open-source real-time models largely outperform offline ones but still trail top proprietary models. Our analysis also reveals that larger model size or higher frame sampling rates do not significantly boost $\mathcal{RTV}\text{-}Bench$ performance, sometimes causing slight decreases. This underscores the need for better model architectures optimized for video stream processing and long sequences to advance real-time video analysis with MLLMs.
Figure 1: Specific examples of different tasks in RTV-Bench.
and
represent the fundamental questions in the question group, and the answers are underlined.
is a dynamically answered question, and the correct answer is determined by the time of the query. The correct answer to the query will change at different times, and we have marked the correct answers that should be responded to at different query times.
1 Introduction
The ability to comprehend and respond to complex real-world scenarios in real-time represents a crucial challenge in the pursuit of general artificial intelligence [6, 24, 10]. Drawing inspiration from the remarkable achievements of large language models (LLMs) in a wide range of tasks [25, 11, 2], the potential of multimodal large language models (MLLMs) is increasingly being realized in the domain of visual scene analysis [48, 41, 5]. The research trajectory of Video-LLMs [21, 27, 4, 35, 33] is progressively advancing from the understanding of short, vision-centric videos [15, 13, 22, 19] to the integration of omni-modal information—including video, audio, and subtitles [3, 7, 17, 32, 18, 26]—for comprehensive long video analysis.
Recently, VStream [45] was the first to attempt to test this capability, with a focus primarily on extending the duration of videos. In addition, both StreamingBench [39] and OVOBench [20] have made varying degrees of improvements in the types and standards of assessment. However, their evaluation of real-time responsiveness is often inadequate, overlooking the capacity to capture transitions and fleeting details from visual input that arrive sequentially, not instantaneously. This limitation highlights the need for a more focused assessment of continuous analysis abilities.
Based on the above considerations, we introduce $\mathcal{RTV}\text{-}Bench$, featuring three core innovations designed to benchmark the continuous analysis capabilities—specifically perception, understanding, and reasoning—of MLLMs within real-time video contexts. First, the Multi-Timestamp Q&A mechanism challenges real-time tracking and state update by posing queries whose answers evolve within a video. Crucially, unlike benchmarks like OVO-Bench [20] that typically introduce different questions at different timestamps, RTV-Bench revisits the same conceptual query, where only the correct answer shifts as the scene unfolds. This approach more rigorously tests the model capacities for continuous analysis in real-time scenarios, surpassing single-query-single-answer and static evaluations. Second, the Hierarchical Question Structure enforces reliable, sequential reasoning by employing a basic-to-advanced design where higher-order questions logically depend on grasping foundational perceptions and understanding, thus mitigating cognitive shortcuts. Finally, Multidimensional Evaluation moves beyond single scores to provide granular diagnostics, assessing performance across eight fine-grained dimensions critical to continuous analysis in dynamic scenarios, offering a more insightful perspective on model capabilities and bottlenecks in handling real-time video information, detailed cases are shown in Figure 1.
Through RTV-Bench’s novel design and the detailed evaluations presented herein, we provide comprehensive insights into the current state of MLLMs for continuous video analysis: Results for all tested models on RTV-Bench reveal significant bottlenecks for existing MLLMs in real-time video analysis, as (1) most models achieve accuracy below 50%. Furthermore, (2) model performance shows minimal correlation with the number of input frames or model size. Notably, (3) models specifically designed for real-time video significantly outperform traditional offline video models, even the lowest-performing real-time model tested (VITA-1.5 [9]) surpasses a traditional offline counterpart (VideoLLaMA2 [4]). Building upon these findings, we further discuss promising research trajectories to enhance MLLM online video processing. We expect that our benchmark, analysis, and outlined future directions will contribute significantly to the development of more robust and capable online video MLLMs.
2 RTV-Bench
Figure 2: Video Categories and Distribution of Question Difficulty and Query Characteristics. (Left) RTV-Bench overs 3 key domains and 16 sub-class video types. (Center) Distribution of question difficulty levels across eight representative task types, measured by percentage-based performance ranges. (Right) Distribution of question queries by video length, categorized into Shallow, Moderate, and Deep levels. The bar heights indicate counts, while the line chart overlays query proportions for each duration bucket.
In this section, we discuss a common challenge in both long video comprehension and real-time video analysis—the continuous analysis capability of MLLMs. To tackle this, we have meticulously curated a benchmark to assess model performance in real-time and continuous scenarios, and developed new metrics to characterize the performance of multimodal large models in this bench.
2.1 Challenge for Real-Time Video Analysis
Existing benchmarks for long video analysis [8, 38] attempt to gauge the true capacity of MLLMs by extending video duration, enhancing the difficulty of QA tasks, and introducing new sub-tasks. Other works [23] focus on real-time video scenarios, evaluating model responsiveness to queries in real-time contexts. However, current models face the risk of memory loss and attention shift, both in long and real-time video contexts. We categorize these challenges as failures in the continuous analysis capability of MLLMs. In video scenarios, MLLMs are considered to possess strong continuous analysis ability only if they can effectively recall prior visual information to accurately respond to user queries (similar to needle-in-a-haystack tasks in natural language settings) and maintain robust perception of incoming streams of new visual data and queries (akin to multimodal multi-turn dialogues).
2.2 Benchmark Overview
The RTV-Bench is designed to assess a model’s ability to perceive, understand, and reason in long video and real-time streaming contexts. Models should be adept at recognizing correct temporal query patterns. For example, during live sports queries (e.g., goalkeeper actions), models must prioritize current contexts over historical data. However, it should also retrieve relevant information from memory, such as identifying the goalkeeper from earlier footage. This necessitates that models not only possess exceptionally high fundamental scene comprehension capabilities, but also demonstrate effective understanding of all stages within the temporal flow—namely, comprehensive spatiotemporal comprehension encompassing the past, present, and future. To this end, we propose a series of sub-tasks for video continuous analysis, addressing both spatiotemporal elements within the video and the intrinsic capabilities of the model. The intrinsic model capabilities are divided into eight categories: Temporal Perception (TP), Scene Perception (SP), Visual Perception (VP), Future Prediction (FP), Phenomenological Understanding (PU), Intent Analysis (IA), Global Understanding (GU), and Spatiotemporal Reasoning (SR), see Figure 1 for details.
2.3 Benchmark Construction
Terminology
The RTV-Bench dataset comprises $N=552$ videos sourced from the internet. A key characteristic is its question structure designed to directly probe temporal dynamics. For each video, questions are organized into sets. Each question set contains a minimum of three multiple-choice questions.
Crucially, the core design principle revolves around time-varying correct answers (Multi-Timestamp QA): within a set of questions, the same underlying conceptual query (for example, ’What is person A holding?’, ’Where is the car heading?’) is effectively evaluated at different points in time within the video. This means that the correct answer option changes depending on the specific timestamp or time interval referenced or implied by the question context or its associated earliest inferable timestamp (Figure 1). This challenges models not just to locate information, but to actively track changes and update their understanding of a situation as it evolves. This design directly targets the assessment of continuous understanding and the model’s sensitivity to temporal changes, see Figure 4 for cases.
The question set features three questions of escalating difficulty. The first two are simpler, while the final question is significantly more complex, demanding integration of broader context or clues. The information and reasoning needed for the final question generally encompass those required for the first two. Successfully answering the complex third question strongly suggests the capability to also answer the simpler ones. This structure evaluates the model’s ability to handle increasing complexity, synthesize comprehensive context, and perform robust analysis.
In addition, we emphasize diversity through the richness of specific sub-scenes and the distribution of video lengths. The RTV-Bench primarily encompasses intelligent driving, sports events, and egocentric videos—categories rich in dynamic information and real-time contextual relevance. Each category includes a variety of sub-categories, as illustrated in Figure 2.
Data Statistic
RTV-Bench comprises 552 videos with a total duration of 167.2 hours (average 18.2 minutes per video) and contains 4,631 QA pairs, for detailed information, refer to Figure  2 and Table  1. The benchmark features diverse question scenarios and evenly distributed video durations, thereby establishing a comprehensive framework for video continuous understanding tasks with rich sub-scenarios and structured problem dimensions.
Video Collection and Filtering
Our video data sources include EgoSchema [28] and publicly available online videos. Unlike most existing benchmarks, we incorporated manual review during the collection phase. Three data collectors sourced videos from various domains and manually excluded highly similar videos, focusing on long videos with high dynamics and real-time needs. These sources ensure targeted and scientific evaluation of video models’ real-time capabilities.
Table 1: Comparison of video QA benchmarks. MT: multi-timestamp questions. Labels: A (automatic), M (manual), A+M (both).
Benchmark MT Num. of QA Avg. Dur. (min) Total Dur. (h) Labels
Video-MME [8] 2.7k 17 254.0 M
MSRVTT [40] 73k 0.25 12.5 A
MSVD-QA [1] 13k 0.17 1.4 A
MovieChat-1k [31] 13k 9.4 156.0 M
MVBench [16] 4k 0.25 34.5 A+M
ActivityNet-QA [43] 58k 1.87 25.0 M
Vstream-Q [45] 3.5k 40 21.0
StreamBench [37] 1.8k 4.5 25.0 M
OVO-Bench [20] 2.8k 6 66.0 A+M
RTV-Bench (Ours) 4.6k 18 167.2 M
Manual Annotation
Rigorous manual annotation by qualified experts was essential for RTV-Bench’s reliability in assessing continuous video analysis. While an LLM (Deepseek [25]) provided initial question templates, human annotators significantly refined questions to target dynamic scenes and temporal reasoning. They meticulously crafted questions where answers change over time and identified the earliest valid timestamp for each answer option in MTQA. This human-centric, multi-annotator process ensured robustness and created evaluations specifically focused on temporal dynamics in video.
Quality Control
To ensure the benchmark’s quality, each video and its corresponding Q&A pairs underwent multiple rounds of review. We manually filtered videos based on length distribution and the presence of sub-scenes examining real-time event changes, resulting in high-quality videos focused on real-time analysis tasks. We conducted manual video-question alignment and precise timestamp checks, utilizing GPT-4 and human review to verify annotation format, translation, and sensitive information.
3 Experiments
3.1 Experiment Setup
Our experiments were conducted across two A800 GPUs to comprehensively evaluate the performance of mainstream MLLMs on our benchmark. The selected models include VideoLLaMA2 [4], VideoLLaMA3 [44], GPT-4O [12], InternLM-XComposer2.5-OmniLive (IXC2.5-OL) [46], VITA1.5 [9], LLaVA-Video [21], LLaVA-OneVision [14], and Qwen2.5-VL [42]. These models were primarily tested with 7B to ensure a fair comparison under similar computational constraints.
Real-Time Video Model vs. Offline Video Model
We compare two model categories: traditional offline models ($M_{\text{offline}}$) and novel real-time online models ($M_{\text{online}}$). These categories differ significantly in architecture, training, and data requirements. Architecturally, $M_{\text{offline}}$ typically employ sequential vision encoder-decoders, often limited by fixed context windows and incurring higher processing latency. In contrast, $M_{\text{online}}$ are designed for continuous, low-latency ingestion of the video stream $V$. They prioritize maintaining an internal state $S_{t}$ that summarizes the video information processed up to time $t$ (i.e., $V[0,t]$), enabling real-time responsiveness. Models like IXC2.5-OL [46] implement this using techniques like modular parallelism and dedicated long-term memory. These architectural distinctions lead to different training paradigms and data needs. $M_{\text{offline}}$ usually rely on end-to-end fine-tuning using standard annotated video datasets. $M_{\text{online}}$, however, often require specialized training strategies (e.g., VITA-1.5’s staged fusion, IXC2.5-OL’s targeted training for memory and interaction) and benefit most from specialized corpora designed for long-duration, interactive streaming scenarios. The ability of $M_{\text{online}}$ to maintain and update state $S_{t}$ is particularly relevant for RTV-Bench’s Multi-Timestamp QA (MTQA) challenge, where the correct answer $A^{*}(Q,t_{q})$ to a query $Q$ depends on the specific query time $t_{q}$. To evaluate both model types on this benchmark, we adapt the testing procedure. For $M_{\text{online}}$, queries $Q$ are presented at their timestamp $t_{q}$, and the models leverage their continuously updated state $S_{t_{q}}$ to generate the answer $A_{M_{\text{online}}}=M_{\text{online}}(Q,t_{q}|S_{t_{q}})$. Since $M_{\text{offline}}$ lack this inherent streaming capability, we simulate real-time interaction for them: when a query $Q_{i}$ is posed at time $t_{q,i}$, we extract and provide only the relevant video segment $V_{i}$ (corresponding to the query’s context). The offline model’s answer is thus based solely on this isolated segment: $A_{M_{\text{offline}},i}=M_{\text{offline}}(Q_{i},V_{i})$.
To evaluate the performance of these models, we employed two metrics: Accuracy and Score.
Accuracy.
The accuracy metric measures the proportion of correct answers provided by the model compared to the ground truth.
Score.
The score metric evaluates the model’s ability to correctly answer advanced-level questions (type q2), contingent upon its demonstrated mastery of prerequisite basic questions (types q0 and q1) within the same question group, thus emphasizing reliable advanced reasoning built upon a solid foundation. Calculation involves a prerequisite check for each group $i$: if all basic questions are correct ($B_{i}=1$), the group contributes points equal to the number of correctly answered q2 questions ($N_{q2,i}^{\text{correct}}$); otherwise ($B_{i}=0$), it contributes zero points. The final score is the ratio of total conditionally awarded points to the total number of q2 questions across all $N$ valid groups (those containing q2). Formula:
$\text{Score}=\frac{\sum_{i=1}^{N}B_{i}\cdot N_{q2,i}^{\text{correct}}}{\sum_{i=1}^{N}N_{q2,i}^{\text{total}}}$
where $N$ is the number of valid groups, $B_{i}$ is the prerequisite indicator (1 if basics are correct, 0 otherwise) for group $i$, $N_{q2,i}^{\text{correct}}$ is the count of correct q2 answers in group $i$, and $N_{q2,i}^{\text{total}}$ is the total count of q2 questions in group $i$. Advantages of this metric include: ensuring foundational accuracy by rewarding advanced correctness only when basics are mastered; reflecting model robustness by penalizing superficial success on complex tasks without fundamental understanding; and aligning with hierarchical learning principles where complex skills build upon simpler ones.
Table 2: Evaluation results on RTV-Bench. Perception, Understanding and Reasoning are different tasks. FQA notes that Foundational Video Question & Answersing without multi-timestamp answers. MTQA notes that Multi-Timestamp Question&Answersing with multi-timestamp answers.
Model #Size Perception Understanding Reasoning FQA MTQA Overall
Acc (%) / Score Acc (%) / Score Acc (%) / Score Acc (%) Acc (%) Acc (%) / Score
Open-Source Offline Video Models
Qwen2.5-VL [42] 7B 33.53/3.53 31.39/3.70 33.65/5.20 31.37 34.37 32.75/4.10
VideoLLaMA2 [4] 7B 40.62/8.67 39.85/7.77 37.49/6.75 45.77 34.95 39.55/7.90
VideoLLaMA3 [44] 7B 37.98/5.83 35.29/5.73 35.78/6.80 38.62 34.91 36.42/6.10
LLaVA-OneVision [14] 7B 35.38/3.97 34.21/4.63 33.57/4.95 35.80 33.58 34.49/4.40
LLaVA-Video [21] 7B 35.83/5.03 33.81/3.77 35.15/5.75 36.28 34.17 34.90/4.80
Open-Source Online Models
VITA-1.5 [9] 7B 45.66/12.80 44.12/11.83 43.37/10.15 55.06 36.32 44.51/11.80
IXC2.5-OL [46] 7B 47.21/15.87 48.22/15.23 46.18/14.45 59.05 38.21 47.33/15.40
Closed-Source Business Models
GPT-4o [12] 51.61/21.90 49.31/20.76 48.71/23.95 56.53 44.73 50.02/22.10
Gemini 2.0 Flash [34] 41.67/11.00 42.71/12.73 41.44/12.05 47.49 38.64 42.00/12.00
3.2 Experiment Results
Online vs. Offline Models.
As shown in Table 2, online models optimized for real-time processing—particularly IXC2.5-OL—surpass offline counterparts in overall performance metrics. IXC2.5-OL achieves 47.33% Accuracy and 15.40 Score, significantly outperforming offline models like VideoLLaMA2 (39.55% Accuracy / 7.90 Score). Furthermore, when comparing online models, IXC2.5-OL demonstrates a clear advantage over VITA-1.5 with improvements of 2.82% in Accuracy and 3.6 points in Score. A notable performance gap emerges in temporal analysis tasks: IXC2.5-OL attains 38.21% Accuracy in Multi-Timestamp Question Answering (MTQA), substantially higher than the 33–35% range typical of offline models. This notable discrepancy suggests that current online models may be promising avenues toward continuous analysis capabilities.
Table 3: Detailed evaluation results on the category of Perception. Temporal Perception (TP), Visual Perception (VP) and Scene Perception (SP).
Method #Size TP VP SP Overall
Acc (%) / Score Acc (%) / Score Acc (%) / Score Acc (%) / Score
Open-Source Offline Video Models
Qwen2.5-VL [42] 7B 32.37 / 2.4 37.48 / 5.8 30.73 / 2.4 33.53 / 3.53
VideoLLaMA2 [4] 7B 39.52 / 7.9 42.49 / 9.4 39.85 / 8.7 40.62 / 8.67
VideoLLaMA3 [44] 7B 37.82 / 6.1 39.24 / 7.6 36.87 / 3.8 37.98 / 5.83
LLaVA-OneVision [14] 7B 35.09 / 3.9 35.86 / 3.8 35.20 / 4.2 35.38 / 3.97
LLaVA-Video [21] 7B 34.07 / 4.8 38.97 / 5.8 34.45 / 4.5 35.83 / 5.03
Open-Source Online Models
VITA-1.5 [9] 46.51 / 12.1 47.09 / 13.2 43.39 / 13.1 45.66 / 12.8
IXC2.5-OL [46] 7B 49.57 / 17.6 49.80 / 16.5 42.27 / 13.5 47.21 / 15.87
Closed-Source Business Models
GPT-4o [12] 48.60 / 18.2 53.59 / 23.4 52.63 / 24.1 51.61 / 21.90
Gemini 2.0 Flash [34] 40.49 / 9.5 45.19 / 16.1 39.34 / 7.4 41.67 / 11.00
Table 4: Detailed evaluation results on the category of Understanding. Phenomenological Understanding (PU), Global Understanding (GU) and Intent Analysis (IA).
Method #Size GU PU IA Overall
Acc (%) / Score Acc (%) / Score Acc (%) / Score Acc (%) / Score
Open-Source Offline Video Models
Qwen2.5-VL [42] 7B 29.11 / 2.2 35.69 / 6.9 29.36 / 2.0 31.39 / 3.70
VideoLLaMA2 [4] 7B 37.34 / 7.6 42.21 / 9.5 40.92 / 6.2 39.85 / 7.77
VideoLLaMA3 [44] 7B 33.54 / 5.8 39.13 / 5.9 33.39 / 4.3 35.35 / 5.33
LLaVA-OneVision [14] 7B 32.07 / 4.3 33.51 / 3.0 37.06 / 6.6 34.21 / 4.63
LLaVA-Video [21] 7B 29.42 / 2.5 35.69 / 3.9 36.33 / 4.9 33.81 / 3.77
Open-Source Online Models
VITA-1.5 [9] 7B 40.30 / 7.2 46.01 / 15.1 46.06 / 13.2 44.12 / 11.83
IXC2.5-OL [46] 7B 43.88 / 11.9 52.17 / 18.7 48.62 / 15.1 48.22 / 15.23
Closed-Source Business Models
GPT-4o [12] 45.02 / 15.7 54.32 / 25.8 48.58 / 20.8 49.31 / 20.76
Gemini 2.0 Flash [34] 35.70 / 10.6 45.65 / 11.3 46.78 / 16.3 42.71 / 12.73
Table 5: Detailed evaluation results on the category of Reasoning. Future Prediction (FP) and Spatiotemporal Reasoning (SR).
Method #Size FP SR Overall
Acc (%) / Score Acc (%) / Score Acc (%) / Score
Open-Source Offline Video Models
Qwen2.5-VL [42] 7B 35.33 / 4.4 33.67 / 4.5 34.50 / 4.45
VideoLLaMA2 [4] 7B 41.47 / 7.5 33.50 / 6.0 37.49 / 6.75
VideoLLaMA3 [44] 7B 38.05 / 6.9 33.84 / 3.9 35.95 / 5.40
LLaVA-OneVision [14] 7B 38.23 / 7.2 28.91 / 2.7 33.57 / 4.95
LLaVA-Video [21] 7B 39.08 / 9.1 31.22 / 2.4 35.15 / 5.75
Open-Source Online Models
VITA-1.5 [9] 7B 47.95 / 12.2 38.78 / 8.1 43.37 / 10.15
IXC2.5-OL [46] 7B 51.88 / 18.1 40.48 / 10.8 46.18 / 14.45
Closed-Source Business Models
GPT-4o [12] 54.67 / 27.1 42.75 / 20.8 48.71 / 23.95
Gemini [34] 2.0 Flash 44.42 / 13.6 38.46 / 10.5 41.44 / 12.05
Open-Source vs. Close-Source Models.
While a performance gap persists compared to leading closed-source models like GPT-4o, state-of-the-art online architectures demonstrate remarkable progress. The online model IXC2.5-OL achieves near-top-tier performance with 47.33% Overall Accuracy and 15.40 Score, substantially outperforming mid-range closed-source systems like Gemini 2.0 Flash. Notably, IXC2.5-OL closes the accuracy gap with GPT-4o to 4.4% in perception tasks(Tables 4) and 1.1% in video understanding, demonstrating competitive performance in multimodal domains. However, limitations emerge in complex reasoning where GPT-4o maintains decisive advantages, especially on complex tasks like Understanding and Reasoning(Tables 4 and 5). This pattern highlights that while modern online models have approached entry-level commercial systems and even challenged premium models in specific competencies, structural innovations remain critical to bridge gaps in advanced cognitive tasks like multi-step reasoning and temporal analysis.
Impact of Model Scales.
To investigate the effect of model scale, we evaluated Qwen2.5-VL across 3B-72B parameters. Figure 3 reveals non-monotonic scaling: medium-sized models (7B-34B) outperform both smaller and larger variants. All metrics fluctuate within a 5% range, with no configuration exceeding 34B’s maximums. This scale saturation implies current model architectures may lack the representational efficiency required for effective real-time processing, emphasizing the need for structural innovations rather than pure parameter growth.
Impact of Frame Numbers.
Figure 3 presents another counter-intuitive finding. While more input frames typically yield better performance in offline evaluations, our real-time setting shows no performance improvement with increased frames. In fact, IXC2.5-OL exhibits significant degradation. This suggests potential limitations in the model’s processing capacity or a breakdown mechanism, such as attention dilution. A potential limitation is that the current frame numbers may not adequately capture all critical moments or key information within long videos. Even with increased sampling density, crucial details might still be missed.
Figure 3: Performance visualization and analysis on RTV-Bench: (a) Visualization of overall Accuracy results; (b) Visualization of overall Score results; (c) Performance impact of varying input frame counts; (d) Performance comparison across different Qwen2.5-VL model scales.
4 Related Work
4.1 Advanced MLLMs towards Real-time Video Analysis
Video-oriented large-scale models represent a highly promising domain with significant real-world potential, giving rise to numerous valuable applications [27, 8]. Video-LLMs have rapidly advanced, moving from analyzing short videos [15, 21] to longer ones [30, 35, 47] using various techniques. However, most research targets offline video analysis. Real-world applications necessitate real-time, continuous perception and reasoning on unfolding video streams. While models like InternLM-XComposer2.5-OmniLive [46], VITA-1.5 [9], and Dispider [29] are exploring real-time capabilities, rigorously evaluating their ability to continuously track and understand dynamic events remains a critical challenge.
4.2 Existing Video Benchmarks
Existing benchmarks primarily focus on offline evaluation using offline videos [8, 36, 38]. These are less suited for assessing how MLLMs track dynamically changing states, as they often use static question-answer pairs. Newer benchmarks tackle real-time and streaming aspects [23, 20], evaluating responsiveness and contextual understanding in online settings. However, RTV-Bench specifically addresses the gap in evaluating continuous perception, understanding, and reasoning. Its key distinction is the use of dynamic question answering, where the correct answer evolves with the video stream, directly probing the MLLM’s ability to maintain and update its understanding of complex, unfolding events over time, complemented by a multi-dimensional evaluation structure.
Figure 4: Comparison of response from different models for the same question on the same video. Green indicates correct answers or timestamps; red indicates incorrect answers or timestamps. This case demonstrates that even current high-performance models struggle to provide high-quality responses (where both the answer and the corresponding timestamp are accurate).
5 Limitations and Future Work
Our benchmark reveals counter-intuitive findings, such as the limited impact of model scale and input frame count on performance, suggesting that MLLM mechanisms for processing continuous video are poorly understood and effective analysis tools are lacking. Furthermore, the current evaluation is primarily limited to the visual modality. Future work will focus on investigating the underlying causes of these phenomena and developing more adapted analytical methods. Concurrently, a key direction involves incorporating important modalities like audio into RTV-Bench to enable a more comprehensive evaluation of continuous perception, understanding, and reasoning in realistic multimodal scenarios.
6 Conclusion
In this work, we introduce RTV-Bench, a Benchmark focuses on evaluating the video continuous analysis capabilities of MLLMs. We proposed relevant evaluation tasks and constructed RTV-Bench, which comprises 552 videos and 4,631 QA pairs. Extensive experiments conducted on diverse models reveal that: Models specifically designed for online video processing exhibit superior video continuous understanding capabilities, whereas simply increasing the number of uniformly sampled frames during training or inference stages currently shows limited effectiveness in enhancing performance on video continuous tasks.
Appendix
The appendix includes the following sections:
  • Section A: Experimental Analysis Supplement.
  • Section B: Framework Design and Application Extensions.
  • Section C: Case.
Appendix A Experimental Analysis Supplement
This section supplements the materials on Score Design and OAE Design, including related content and partial experimental data.
Score VS Accuracy
In Section 3.1, to comprehensively evaluate the performance of these models, we design two metrics: Accuracy for tack-specific correctness and Score for dynamic reasoning consistency. The Score metric is crucial for revealing reliable, hierarchical reasoning beyond simple Accuracy. While GPT-4o’s Accuracy gain over IXC2.5-OL is moderate (50.02% vs. 47.33%), its substantially higher Score (22.10 vs. 15.40) indicates a more robust reasoning process, reliably building upon foundational understanding to address complex queries. Conversely, lower Scores relative to Accuracy, common among open-source models, suggest instability in multi-step reasoning. Thus, as shown in Figure 3, the Score metric effectively quantifies the deeper, more reliable comprehension capabilities demonstrated by models like GPT-4o on RTV-Bench’s challenging tasks and it can distinguish model performance more clearly.
With the aim of validating the discriminative power of the Score metric, we conducted statistical analysis on 1,527 multi-timespan QA-triples, as shown in A.1. The result revealed a significantly positive correlation between foundational reasoning(Q1-Q2) and subsequent conplex reasoning performance(Q3). This demonstrates that models achieving higher accuracy on elementary visual perception tasks exhibit proportionally stronger performance on advanced temporal reasoning tasks.
Table A.1: Accuracy Distribution of Question 3 Conditioned on Preceding Question Performance: Q3 Accuracy when at least one of the preceding questions (Q1 or Q2) was answered correctly, versus Q3 accuracy when both preceding questions (Q1 and Q2) were answered incorrectly.
[…]
GPT-4o [12] 297 64
IXC2.5-OL [46] 7B 499 81
LLaVA-OneVision [14] 7B 309 182
LLaVA-Video [21] 7B 308 195
Such alignment shows our core design philosophy for RTV-Bench: hierarchical reasoning capabilities depend on robust foundational perception, mirroring human cognitive processes in real-time video understanding. Furthermore, it also illustrates the rationality of Score as a real-time video metric, which addresses critical limitations of conventional single-answer accuracy metrics in assessing continuous reasoning dynamics.
OAE Design Overview
Table A.2: Detailed evaluation results on the category of OAE. Object, Action and Event.
Method #Size Object Action Event
Acc (%) / Score Acc (%) / Score Acc (%) / Score
Open-Source Offline Video Models
Qwen2.5-VL [42] 7B 31.38 / 2.8 34.75 / 4.9 32.63 / 4.5
VideoLLaMA2 [4] 7B 40.39 / 8.2 40.25 / 8.6 38.69 / 6.8
VideoLLaMA3 [44] 7B 34.31 / 4.5 37.77 / 7.4 34.21 / 3.9
LLaVA-OneVision [14] 7B 34.31 / 5.0 35.82 / 4.9 33.42 / 3.4
LLaVA-Video [21] 7B 36.10 / 6.1 35.40 / 4.7 33.88 / 3.8
Open-Source Online Models
VITA-1.5 [9] 7B 47.39 / 13.7 44.09 / 11.6 42.85 / 10.3
IXC2.5-OL [46] 7B 49.89 / 17.2 46.34 / 14.6 46.61 / 15.4
Closed-Source Business Models
GPT-4o [12] 50.63 / 23 50.97 / 22.3 49.01 / 20.9
Gemini [34] 2.0 Flash 42.66 / 12.2 43.34 / 11.5 40.24 / 12.4
During evaluation, we also incorporated the Object-Action-Event framework( Table B.1) as part of our analytical scope, designed to assess video comprehension from multiple agent-centric perspectives. For example, in live sports scenarios, we evaluate three perspectives: objects (e.g., players appearing or disappearing during an offensive play), actions (e.g., dynamic maneuvers by offensive players), and events (e.g., real-time offensive strategies deployed in the midfield).
OAE Accuracy and Score Analysis
In Table B.1, GPT-4o maintains the leading positions in all three dimensions of the OAE , and all online models demonstrate significantly higher accuracy and scores compared to offline models, particularly in the Object dimension, this highlights that the perspective design of OAE requires exceptionally strong capabilities in continuous analysis tasks. Overall, the models show no significant gaps in accuracy and scores across the three dimensions. Most offline models achieve their best performance in the Action aspect, while online models excel particularly in Object. Notably, nearly all models exhibit the lowest accuracy and scores in the Event dimension, indicating that continuous analysis tasks with higher complexity remain a formidable challenge for current systems.
Table A.3: Analytical Object Taxonomy: RTV-Bench systematically formulates an object-action-event framework, with explicit definitions of core characteristics and discriminative criteria for each category.
Category Dimensions
Spatiotemporal Elements Objects: Physical entities appearing in video frames.
Actions: Dynamic behaviors performed by objects.
Events: Complex occurrences combining objects and actions.
Appendix B Framework Design and Application Extensions
This section supplements the materials on the rationale for and necessity of evaluation dimension design, while providing an extended analysis on the broader utility of the dataset.
Methodological Foundations for Assessing Continuous Analysis Capabilities
To further elaborate on video continuous analysis capabilities introduced In Section 2.2, we formalize the foundational definitions and evaluation protocols for this capacity. Primarily, models are required to possess perception, understanding, and reasoning abilities comparable to state-of-the-art offline video models before addressing real-time streaming contexts. This requirement motivates our two-stage QA design, as elementary offline video analysis capabilities intuitively form the prerequisite for advanced temporal reasoning.
Furthermore, models must demonstrate proficiency in recognizing correct temporal query patterns due to three critical demands inherent to real-time applications: enhanced perception capabilities in highly dynamic scenarios requiring rapid and precise visual processing, deepened understanding of ongoing events under temporal continuity constraints, and effective reasoning about future trajectories based on evolving contextual cues.
Why evaluate across perception, understanding, and reasoning dimensions?
To elucidate the necessity of three-dimensional categorization, we subsequently analyze perception, understanding, and reasoning respectively.In egocentric driving environments with high-dynamic scenarios, the video continuous analysis capability fundamentally addresses two aspects of persistent navigation: 1) holistic operational state awareness (e.g., current traffic condition assessment, historical route context) through the integration of offline and online video processing capabilities, and 2) perception speed, comprehension, and analytical capabilities for sudden real-time events.
Regarding perception design, this corresponds to detecting abrupt environmental changes during navigation, such as traffic light transitions, emergent vehicles, and pedestrians - scenarios where conventional video models exhibit critical deficiencies in temporal responsiveness. As visualized in Figure 4, existing architectures struggle to adapt to real-time variations in high-dynamic settings. We systematically decompose this capability into three sub-dimensions: Temporal Perception, Scene Perception and Visual Perception.
Regarding understanding design, it addresses the interpretative capacity for sudden operational changes during driving. For the understanding design, it pertains to scenarios involving the comprehension of abrupt changes during driving, such as interpreting traffic signal indications, road sign semantics, and the rationale behind preceding vehicles’ maneuvering strategies. Previous models have been shown to fall short of fundamental requirements in both processing speed and interpretative depth. For instance, during traffic signal transitions or lane-changing events, current models struggle to detect such changes within reasonable timeframes. In complex real-time traffic scenarios, beyond insufficient processing speed, existing systems also largely fail to meet advanced comprehension requirements for situational awareness. Through systematic categorization and abstraction of diverse scenarios, we decompose this dimension into three sub-components: Intent Analysis, Phenomenological Understanding and Global Understanding.
Regarding reasoning design, it encompasses scenarios requiring continuous temporal analysis, such as traffic signal duration estimation, and anticipating preceding vehicles’ strategies, these demand sophisticated continuous analysis capabilities. Based on the dual requirements of historical context integration and prospective forecasting, we architect this dimension into two sub-dimensions: Spatiotemporal Reasoning and Future Prediction.
Table B.1: Core Evaluation Dimensions: RTV-Bench systematically defines eight essential evaluation dimensions for continuous video understanding systems, accompanied by formal characterizations and discriminative criteria for each dimension.
Category Dimensions
Model Capabilities Temporal Perception (TP): Recognizing temporal sequence and duration.
Scene Perception (SP): Understanding holistic environment and layout.
Visual Perception (VP): Detecting fine-grained visual features.
Future Prediction (FP): Anticipating future developments.
Phenomenological Understanding (PU): Interpreting surface phenomena.
Intent Analysis (IA): Inferring actor motivations.
Global Understanding (GU): Grasping video context.
Spatiotemporal Reasoning (SR): Logical deduction from observations.
Why cross-apply the OAE design with eight evaluation dimensions?
We observe notable limitations in the dimensional design frameworks of current mainstream benchmarks, suggesting areas that warrant systematic refinement. Current evaluation taxonomies frequently include components such as Object Recognition and Action Reasoning, yet conspicuously omit complementary dimensions like Object Reasoning and Action Recognition. This prevalent pattern reveals significant arbitrariness in dimension partitioning, resulting in evaluation frameworks whose systematic rigor and scientific credibility remain fundamentally compromised. To address these limitations, we propose a novel taxonomy that first independently categorizes analytical subjects and evaluation dimensions, subsequently implementing cross-categorization to establish a structured and systematic grid framework for dimensional organization.
Figure B.1: Our dataset construction pipeline. We develop a dataset generation pipeline consisting of three stages to create RTV-Bench: Video Collection, Templates Design and Human Annotation.
Supplementary Analysis of Dataset Value
This work introduces the Real-time Video Understanding Benchmark (RTV-Bench), which holds significant potential to advance video comprehension research and its applications. By defining core capabilities required for real-time video processing through high-dynamic video evaluation, RTV-Bench establishes continuous analysis capabilities as critical developmental objectives. The benchmark demonstrates a novel paradigm for online video capability assessment, potentially accelerating the enhancement of core functionalities in online video understanding models and expediting their deployment in real-time scenarios such as embodied AI, autonomous driving, and live sports commentary.
Experimental analyses using RTV-Bench reveal substantial deficiencies in existing video models’ continuous analysis capabilities while illuminating potential improvement pathways. Our results indicate that increasing model scale or video sampling frequency yields diminishing returns. The substantial performance gap between online and offline models in continuous analysis capabilities not only validates RTV-Bench’s design efficacy but also suggests architectural innovation as a viable development direction.
The RTV-Bench framework is constructed upon three foundational principles: 1) Balanced Taxonomy Design: Prioritizing structural rationality in evaluation dimensions while implementing novel task formulations. 2) QA Integrity Assurance: Maintaining high-quality question-answer pairs while mitigating validity threats from oversimplified QA designs, particularly accuracy inflation through correct guessing in multiple-choice formats. 3) Capability Anchoring: Conducting specialized evaluations for real-time video processing while systematically monitoring potential degradation in foundational video understanding competencies. Furthermore, the rigorous RTV-Bench generation pipeline B.1 also ensures high-quality QA pair production B.2.
Figure B.2: This figure presents a word cloud visualization derived from QA question corpora, explicitly highlighting the most frequently occurring terms through typographic prominence.
Broader Impact
By focusing on real-time video comprehension, RTV-Bench rigorously evaluates models’ continuous temporal modeling and dynamic scenario comprehension capabilities. This benchmarking framework directly addresses critical gaps in deploying video understanding systems for latency-sensitive applications, including embodied AI navigation requiring sub-second environmental feedback, autonomous driving systems dependent on frame-by-frame hazard anticipation, and live sports commentary generation demanding real-time event contextualization.
Regarding potential negative societal impact, while RTV-Bench can better facilitate the development and application of real-time video tasks, these technologies may concurrently promote the creation of more sophisticated real-time surveillance systems, which inherently raises privacy concerns. Therefore, the advancement of related technologies must be conducted under the supervision of robust privacy protection measures and ethical standards.
Appendix C Case
We present additional examples from RTV-Bench, covering eight evaluation dimensions and three analytical perspectives, to demonstrate the holistic and systematic framework for assessing continuous analysis capabilities.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly reflect the paper’s core contributions and research scope. Specifically, they explicitly present the research motivation for RTV-Bench, identify the limitations of existing paradigms, and introduce continuous analysis capabilities along with the associated evaluation protocol.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of the work are detailed in Section 5
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: As a dataset work, this paper does not contain theoretical derivations.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The experimental section of this paper includes dataset evaluation benchmarks, with the dataset open-sourced on public hosting platforms and accompanied by a Croissant metadata file submission.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: RTV-Bench has been open-sourced on public hosting platforms and submitted with a Croissant metadata file.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental configurations for evaluation are comprehensively detailed in Section 3.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We do not report error bars or statistical signiffcance tests in this paper due to the extensive computational cost associated with our experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The experimental configurations for evaluation are comprehensively detailed in Section 3.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in this paper fully conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The broader impacts of this work are detailed in Section B
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This work does not contain high-risk content.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All referenced papers are fully cited with complete bibliographic entries in the References section.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [No]
  64. Justification: The documentation has not yet been finalized and made publicly available in public repositories as of the current stage of development.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve research with human subjects or crowdsourcing.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [No]
  79. Justification: This work is a dataset project where LLMs are solely used as evaluation targets and utilized to inform the dataset construction pipeline.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.